# Sound Source Localization for Spatial Mapping of Surgical Actions in Dynamic Scenes

### v1

@@ 5 Discussion
The event detection module showed promising results, especially for chiseling and sawing, where events could be detected within a maximum offset of one frame for chiseling (corresponding to at most 20ms20\>ms), and a maximum offset of three frames (corresponding to at most 60ms60\>ms) with high accuracy. Especially, for sharp burst events, such as chiseling, an accurate detection is crucial. For sawing, the recordings contained only one onset and then a continuing sawing action, as visible in the waveform plot in Figure 3 which results in only one event present per sequence. Therefore, a relaxed criterion for the detection of the onset is tolerable. The detection of sawing events presented a more challenging problem, where we tolerate a maximum offset of 10 frames for a detected event (corresponding to at most 200ms200\>ms) and obtain decreased performance compared to the other surgical actions that have been analyzed within this work. This decreased performance can mainly be attributed to the fact that the surgeons turned on the device in the beginning of the sequence and kept drilling into the plastic bone while the power tool was running, resulting in only very subtle sound differences between drilling in the air and drilling into the rather soft plastic bone. To mitigate this issue, beam steering or noise canceling could improve the results, which is, however, beyond the scope of the presented work.
@@
The event localization module proved that the fusion of audio-visual information is an effective approach to detect regions of interest with minimal computational overhead. While this method can directly be employed to inform down-stream tasks, such as scene graph generation or instrument tracking, we also see potential as an initial filtering stage that allows more computationally intensive methods to focus solely on the regions of interest. Despite the simplicity of our approach, $77\%$ of the events were successfully localized with a 3D bounding box intersection over union of at least $0.1$. This accuracy is enough to generate semantic scene states, i.e., scene graphs for surgical context understanding. Due to the low spatial resolution of the acoustic camera the localization accuracy is limited, as indicated by the steep decrease in recall under stricter intersection over union thresholds. Note however that our predictions were compared to posed ground-truth bounding boxes, whereas our approach does not explicitly optimize the rotation of the estimated bounding boxes. We argue that for region of interest detection a coarse localization is an acceptable trade-off for low latency. For tasks with higher localization requirements, the accuracy could be improved by occlusion-aware fusion of acoustic features into the scene representation. Future extensions using beam-steering based on the 3D scene geometry and a projection of more informative acoustic features, for example in the form of neural implicit sound representations, could bring further enhancements.
@@
The exploratory character of our experiments inherently comes with some limitations. While the presented approach is real-time capable in theory, our current implementation only supports offline processing due to limitations in the software and API of the acoustic camera. We employ a single-view point cloud as the dynamic 3D visual representation of the surgical scene, which is the basis for mapping the localized acoustic activity. While the ZED2i RGB-D camera produces high-quality point clouds, the preservation of fine details and completeness of the scene is limited. To obtain a higher-quality visual representation, future research should utilize multi-view reconstructions to reduce occlusions and investigate alternative scene representations such as dynamic gaussian splats.
@@ 6 Conclusion
Acknowledgement
This work has been funded by the Swiss National Foundation under the funding scheme ”Spark” with grant number 228813. The experiments were supported by the OR-X - a Swiss national research infrastructure for translational surgery and associated funding by the University Hospital Balgrist. Imaging was performed with support of the Swiss Center for Musculoskeletal Imaging (SCMI). We furthermore thank Dr. Matthias Brechbühl from Norsonic Brechbühl AG and Carsten Hessenius, Andy Meyer, and Michael Markus Ackermann from GFaI e.V. and gfai tech GmbH for the support that made this work possible.

### v2

@@ 5 Discussion
The event detection module showed promising results, especially for chiseling and sawing, where events could be detected within a maximum offset of one frame for chiseling (corresponding to at most 20ms20\>ms). Especially, for sharp burst events, such as chiseling, an accurate detection is crucial. For drilling and sawing, the actions span over a longer period of time, as visible in the waveform plot in Figure 3. Therefore, a relaxed criterion for the detection of the onset is tolerable. The detection of drilling events presented a more challenging problem which can mainly be attributed to the fact that the surgeons turned on the device in the beginning of the sequence and kept drilling into the plastic bone while the power tool was running, resulting in only very subtle sound differences between drilling in the air and drilling into the rather soft plastic bone. To mitigate this issue, beam steering or noise canceling could improve the results, which is, however, beyond the scope of the presented work.
@@
The event localization module proved that the fusion of audio-visual information is an effective approach to detect regions of interest with minimal computational overhead. While this method can directly be employed to inform down-stream tasks, such as scene graph generation or instrument tracking, we also see potential as an initial filtering stage that allows more computationally intensive methods to focus solely on the regions of interest. Despite the simplicity of our approach, $84\%$ of the events were successfully localized with a 3D bounding box intersection over union of at least $0.1$. This accuracy is enough to generate semantic scene states, i.e., scene graphs for surgical context understanding. Still, the proposed learning-free baseline and the low resolution of the acoustic camera limit the localization accuracy, as indicated by the steep decrease in recall under stricter intersection over union thresholds. For tasks with higher localization requirements, our localization baseline could be replaced with a deep learning-based approach to enable the extraction of more descriptive multi-modal features, which may enable a more precise sound source localization. An additional potential direction for future research is the projection of more informative acoustic features by evaluating advanced beamforming algorithms. In this context, our current implementation uses a standard delay-and-sum beamforming algorithm that is part of the commercial software shipped with the acoustic camera used within this study.
Our implementation has a latency of about 250ms250\>ms for the event detection and about 100ms100\>ms for the localization stage (measured on a downsampled point cloud with $\frac{1}{64}$ of the original size). Additional optimizations such as GPU-based acceleration and precompiled methods could further decrease these results. While we cannot measure the latency of the beam-forming algorithm due the absence of an API in the closed-source software, comparable methods take around 200ms200\>ms [jansen2025delay]. Our current implementation only supports offline processing for the same reason. Future works should address these limitations.
The exploratory character of our experiments inherently comes with some limitations. The presented dataset is limited in scale and diversity, containing sequences of three representative surgical action, and does not represent the full complexity of a real surgery. Still, our preliminary results show the potential of our novel multimodal audio-visual surgical scene representation and motivate further research to investigate various aspects for improvement, e.g., a wider range of surgical activity, advanced beamforming algorithms, high-quality 3D scene representations and their enrichment via the integration of additional low-level perception algorithms. During the experiment, we used a plastic bone and foam soft tissue model that was moved frequently to introduce variation. The robustness of the proposed learning-free localization approach should be subject to future studies that should evaluate realistic patient models, as well as obtaining larger amounts of data with more variation.
@@
We employ a single-view point cloud as the dynamic 3D visual representation of the surgical scene, which is the basis for mapping the localized acoustic activity. While the ZED2i RGB-D camera produces high-quality point clouds, the preservation of fine details and completeness of the scene is limited. To obtain a higher-quality visual representation, future research should utilize multi-view reconstructions to reduce occlusions and investigate alternative scene representations such as dynamic gaussian splats.
@@ 6 Conclusion
Declarations
Acknowledgements This work has been funded by the Swiss National Foundation under the funding scheme ”Spark” with grant number 228813. The experiments were supported by the OR-X - a Swiss national research infrastructure for translational surgery and associated funding by the University Hospital Balgrist. Imaging was performed with support of the Swiss Center for Musculoskeletal Imaging (SCMI). We thank Dr. Matthias Brechbühl from Norsonic Brechbühl AG and Carsten Hessenius, Andy Meyer, and Michael Markus Ackermann from GFaI e.V. and gfai tech GmbH for their generous support.
Competing Interests The authors have no relevant financial or non-financial interests to disclose.
Ethics Approval Since this study did not involve human or animal subjects, ethics approval is not applicable.
Informed Consent As this research did not involve human participants, informed consent is not applicable.